# Learning to Detect and Segment Mobile Objects from Unlabeled Videos

### v1

@@ Abstract
Embodied agents must detect and localize objects of interest, e.g. traffic participants for self-driving cars. Supervision in the form of bounding boxes for this task is extremely expensive. As such, prior work has looked at unsupervised object segmentation, but in the absence of annotated boxes, it is unclear how pixels must be grouped into objects and which objects are of interest. This results in over-/under-segmentation and irrelevant objects. Inspired both by the human visual system and by practical applications, we posit that the key missing cue is motion: objects of interest are typically mobile objects. We propose MOD -UV, a Mobile Object Detector learned from Unlabeled Videos only. We begin with pseudo-labels derived from motion segmentation, but introduce a novel training paradigm to progressively discover small objects and static-but-mobile objects that are missed by motion segmentation. As a result, though only learned from unlabeled videos, MOD -UV can detect and segment mobile objects from a single static image. Empirically, we achieve state-of-the-art performance in unsupervised mobile object detection on Waymo Open [44], nuScenes [4], and KITTI [17] Dataset without using any external data or supervised models. Code is publicly available at %definitionshttps://github.com/YihongSun/MOD-UV.
@@ 1 Introduction
Embodied agents such as self-driving cars must detect and localize objects of interest such as traffic participants to operate safely and effectively. Today, building such a detector requires the expensive and laborious annotation of millions of boxes over thousands of images. This process is so expensive that the largest detection dataset is orders of magnitude smaller than classification datasets and has much fewer classes. The limited set of classes further runs the risk of missing important object categories (e.g. snowplows for self-driving cars).
@@
These concerns have motivated research into unsupervised object detection techniques that automatically discover objects from unlabeled data [51, 52, 6]. Under the hood, these techniques use self-supervised features to segment unlabeled images and produce candidate object annotations which are then used to train a detector. However, while promising, these approaches often produce many uninteresting and irrelevant “objects” in cluttered scenes (e.g. roads) and over- or under-segment objects of interest (e.g. multiple detections partitioning a large bus and a single detection grouping a row of parked cars). These failures shouldn’t be surprising: after all, how can a completely unsupervised feature representation encode that one assortment of windows, doors and wheels belongs together as an object of interest, but is a separate object from other nearby windows, doors and wheels?
In this paper, we argue that a key missing cue is motion. As humans, we know to group the windows, doors and wheels together because we have seen this assortment move together, also known as Gestalt principle of common fate [34]. Having seen this assortment move together multiple times in the past, we can now group these correctly into an object even when the object (a car) is static. Indeed, motion-based grouping is one of the first forms of grouping to appear developmentally in human infants [43] and can potentially bootstrap other grouping cues [33]. Motion is important not just perceptually, but also practically: Robots performing navigation tasks must plan their trajectories carefully around objects that can move. Identifying mobile objects is also important in manipulation tasks because it provides a view of the available affordances.
@@
With these motivations in mind, we propose an unsupervised approach to build a detector for mobile objects as shown in Figure 1. Our approach trains on unlabeled videos alone and produces a mobile object detector that can run on static frames. We build upon recent advances in unsupervised learning of depth and motion segmentation from unlabeled video to automatically produce pseudo training labels without supervision. However, such unsupervised motion segmentation poses two key challenges. One, it only identifies moving objects and does not include objects (e.g. parked cars) that are static but mobile. Two, it only identifies nearby objects, since the pixel motion of faraway objects is too subtle to discern. We propose a new training scheme to address these challenges, resulting in a final mobile object detector that detects more mobile objects than the initial motion segmentation.
@@
We argue that motion is an important cue for learning what makes an object.
@@
We propose a new training scheme that trains on unlabeled video to produce a mobile object detector that can run on static images.
@@ 2 Related Work
Learning to identify and localize objects from unlabeled images is a challenging task, as object information must be obtained without any explicit human annotations. A long line of work seeks to discover prominent objects in large image collections. Early work used probabilistic Hough transforms to match proposals across images[9]. This approach was improved upon by framing the problem as an optimization problem over a graph where nodes are object proposals and edges are visual similarities[48] and scaled up to larger collections [49, 50]. However, this approach is fundamentally limited by the quality of the object proposals. More recently, Locatello et al. [30] and DINOSAUR [36] consider object discovery as object-centric learning [18] and leverage Slot Attention to decompose a complex scene into independent objects. However the reconstruction objective is difficult to scale and is likely to discover irrelevant objects as well.
More recent work has relied on the fact that bottom-up segmentation algorithms when applied to self-supervised pretrained representations yield good object proposals. LOST [39], FOUND [40] and FreeSOLO [51] generate pseudo mask labels from DINO [7] features to train their downstream detectors. Similarly, MaskDistill [47] distills a single class-agnostic initial mask from the affinity graph produced by DINO [7]. TokenCut [53] and CutLER [52] use Normalized Cuts [37] instead to generate pseudo-labels. HASSOD [6] leverages hierarchical adaptive clustering to generate pseudo-labels, which improves the detection of small objects and object parts. This line of work now produces detectors that can run on static images, similar to our work. However, the detected objects can often be irrelevant (e.g. buildings and road) or over- or under-segment objects of interest. In contrast, our proposed MOD -UV discovers and detects a more meaningful and practical set of mobile objects instead, and can learn from their apparent motion in unlabeled videos only without relying on any additional datasets.
@@
In addition to unlabeled images, 3D information can also be leveraged for discovering objects. Herbst et al. [21] reconstructs the same scene from multiple traversals via RGB-D inputs and identify discrepancies as discovered mobile objects. MODEST [57] extends upon this and builds an unsupervised 3D detector from traversals of LiDAR scans. Garcia et al. [16] discovers salient objects by late-fusing color and depth segmentation from RGB-D inputs. Tian et al. [46] generates candidate segments via 3D point cloud segmentation from LiDAR recordings. In comparison, MOD -UV does not require any additional sensors or modalities beyond unlabeled videos, which allows our method to work in more general settings.
2.0.3 Unsupervised Video Segmentation.
Another class of related work segments out the moving objects from a video without annotations. Here the input is a video and the output is a binary segmentation of the moving objects. Kwak et al. [25], Xie et al. [54], Lamdouar et al. [26], Yang et al. [56], Zhang et al. [58], Zhou et al. [59], and FODVid [41] perform this task by leveraging optical flow information from an input video. Additionally, Lian et al. [28] improves upon cases of articulated/deformable objects and shadow/reflections by relaxing the common fate assumption. Another line of work uses a reconstruction objective to identify the moving object. Du et al. [13] models explicit physical dynamics and reconstructs 3D object geometry by exploiting motion cues in videos. Bao et al. [2] improves training for object-centric representation via an additional motion segmentation regularization, while SAVi++ [14] incorporates depth values from LiDAR sensor when training an object-centric video model. Aydemir et al. [1] extends DINOSAUR [36] to video learning while proposing a masking strategy for tokens in the feature space for efficiency and regularization. Unlike MOD -UV, these approaches do not build a static image detector. However, the output segmentation can be used as an initialization for our approach.
Closer to our work, Croitoru et al. [11] and Choudhury et al. [10] train a single-frame binary segmentation network on frames of a video as input and leverage object motion as supervision. Furthermore, LOCATE [42] applies graph-cut to obtain binary motion mask from DINO [7] and optical flow feature similarities, which in turn is treated as pseudo-labels for bootstrapped self-training of a downstream segmentation network. However, these techniques can only detect a single salient object per frame. In contrast, MOD -UV generalizes to multi-object detection during inference beyond single-object saliency detection.
@@ 3 Method
Problem setup: We assume an uncurated collection of unlabeled videos as input. In particular, we assume that these videos are obtained by an embodied agent observing, and optionally acting in the world. Solely from the unlabeled videos, the goal is to learn a detector that operates from a single frame and can detect and segment all mobile objects that can move of their own volition.
@@ 3.1 Initialization with Unsupervised Motion Segmentation
A key insights in MOD -UV is that if an object can move, it is likely that it does move many times in the collected data. Thus, we start by identifying moving objects in the videos; they can be initial seeds for learning about mobile objects.
@@
As shown in the left of Figure 2, the initial pseudo-labels , while having high precision, fail to capture the large static objects, e.g. the black sedan in the bottom. However, a parked black sedan looks the same as a moving black sedan if all one has is a single frame. In other words, moving and static objects are indistinguishable when observed from a single frame.
@@
Also, since came from a heavily-regularized detector, we maintain and finetune the pseudo-labels for larger objects by training a second detector from scratch in parallel, on the training pair without down-scaling or padding. Notably, this is different from traditional scale jittering, since the singular detector would be discouraged from detecting small objects at larger scales due to the limitations in . Upon convergence, we have a Large-object detector trained at original scale and another Small-object detector trained at a reduced scale. After aggregating their predictions and resolving conflicting proposals222Details on the mask aggregation are found in Appendix 0.A.1., we obtain the final pseudo-labels, . We note that separating out large and small object detectors in this way has been explored in supervised face detection [22].
@@ 3.3 Implementation details
We follow the official code release by Dynamo-Depth[45] and train the system on Waymo Open [44] only to obtain the motion segmentation network and monocular depth estimator. In the initial pseudo-label generation stage, we binarize the estimated motion mask via a threshold of and cluster the pseudo 3D points via DBSCAN [15] using a 10-by-10 local pixel neighborhood connectivity, Eps at 0.03, and required number of samples for core point at 100.
@@
During Moving2Mobile, we train a detector for 3 epochs, with scale jittering from to . Since early-stopping is applied, we adopt a lower confidence threshold of to compute the next round pseudo-labels . During Large2Small, we train both the large and small detectors for 20 epochs, with fixed scale jittering at and , respectively. As both are trained to convergence, we adopt a higher confidence threshold of and , respectively, to compute the next round pseudo-labels with aggregation. For Final round, we train the detector from scratch for 20 epochs, with scale jittering from to . The self-training in MOD -UV takes 27 hours on 1 NVIDIA A6000 GPU.
@@ 4.1 Experimental Setup
4.1.1 Datasets.
For evaluation, we focus our attention to self-driving datasets since uncurated, unlabeled video data is available and detecting mobile objects is of interest for autonomous vehicles. We train both MOD -UV‡ and MOD -UV on Waymo [44] and compare our performance with baselines on Waymo. We then also evaluate generalization to nuScenes [4], KITTI [17], and COCO [29].
We split the 798 sequences from Waymo train set into 762 for training and 36 for validation. After method development concludes, we evaluate on the held-out 1,881 test images (averaging 28.4 mobile instances per image) from Waymo val set [32]. Additionally, nuScenes, KITTI, and COCO are only used for evaluating generalization. We test on 3,249 front-camera images (average of 8.2 mobile instances per image) from the nuImage validation set for nuScene and 7,481 images (average of 6.9 mobile instances per image) in the 2D Detection training set for KITTI. For COCO, we evaluate on 870 images (average of 3.8 mobile instances per image) in COCO val 2017 that contain ground vehicles.333Additional dataset information is found in Appendix 0.B.
4.1.2 Baselines.
For the task of unsupervised mobile object detection, there is no directly comparable baselines to the best of our knowledge. Therefore, we consider methods for unsupervised object detection, namely CutLER [52] and HASSOD [6], as the closest points of comparison.
CutLER [52] uses normalized cuts on DINO features (trained on ImageNet) to generate pseudo labels that are used to train a Cascade Mask R-CNN [5] with ResNet-50 backbone. We evaluate the official checkpoint.
HASSOD [6] is a follow-up to CutLER. It discovers objects on COCO [29] via a hierarchical adaptive clustering of DINO features (trained from ImageNet). The hierarchical clustering yields three “levels”: objects, object parts, and object sub-parts. For consistency, we consider all three hierarchical levels for evaluation. As before, we use its official released checkpoint. Since both MOD -UV‡ and MOD -UV are trained on Waymo [44], we also consider a version of HASSOD solely trained on Waymo Open [44], which we denote as HASSOD†.
CutLER[52] and HASSOD[6] are trained to detect all objects in an image regardless of their ability to move. However, our evaluation and approach is focused on mobile objects. We therefore also consider an “oracle” version of these baselines where we additionally remove any CutLER and HASSOD predictions that overlap by less than 0.1 in IoU with the ground truth instances. These oracles, namely CutLER∗, HASSOD∗, and HASSOD†∗, are grayed out in the tables to indicate additional ground-truth-based filtering.
@@
4.1.3 Metrics.
We evaluate both Average Recall (AR and AR) and Average Precision (AP and AP). Since the task is unsupervised and no semantic information is given during training, we follow prior arts [52, 6] and evaluate class-agnostic AR and AP by treating all predicted and ground truth instances as a single class of “foreground” or “mobile” objects.444Complete benchmarks, repeated training performance, and additional ablation studies are found in Appendix 0.C.
@@ 4.2 Unsupervised Mobile Object Detection and Segmentation
Recall. We report AR0.5 (average recall with IoU), AR, AR, AR, and AR for both box and mask predictions. MOD -UV significantly outperforms prior arts across all recall metrics except the recall for large objects where it is comparable. The improvement is especially large for small objects (4.7 higher AR than the nearest competitor). Compared to a supervised Mask R-CNN trained on COCO [29], MOD -UV closes the gap in Box AR from 11.6 to 4.8. Our gains are also much larger on the AR0.5 metric (nearly 2 prior state-of-the-art). This suggests that we detect significantly more objects than prior work, but their localization can be improved. Even so, on overall AR we still show a -point improvement. Additionally, although HASSOD† is trained in-domain on Waymo, it underperforms HASSOD on Waymo, which we suspect is due to the uncurated nature of self-driving scenes.
Precision. We report AP at an overlap threshold of 0.5, as well as AP, AP, AP and AP. Since CutLER and HASSOD are trained to detect all objects in an image regardless of their ability to move, we also compare to oracle versions of these techniques with ground-truth-based filtering. On Waymo, MOD -UV significantly outperforms prior arts across all precision metrics. Even with ground-truth filtering (in gray), MOD -UV still consistently outperforms baselines on all precision metrics (except for large objects where it is comparable). Specifically, MOD -UV outperforms the nearest competitor (with ground-truth filtering )by 4.1 on Box AP and 3.4 on Mask AP, and is 4.5 higher on AP. Intriguingly, compared to a supervised Mask R-CNN trained on COCO [29], MOD -UV closes the gap in Mask AP from 4.3 to 0.8 points.
@@
Recall. As shown in Table 2 and 3, on nuScenes and KITTI, MOD -UV consistently outperforms prior arts across all AR metrics except for large objects, achieving a more than 1.5 improvement on AR0.5 over the nearest competitor on nuScenes. MOD -UV also shows large gains on small objects, improving AR by 2.4 on nuScenes and over 1.7 on KITTI.
Finally, we evaluate on COCO, which is in-domain for HASSOD and a big domain shift for MOD -UV that is trained on Waymo. Notably, MOD -UV maintains superiority on AR and AR at IoU, while being comparable to HASSOD on AR and AR.
Precision. This improvement is also seen in AP. On both nuScenes and KITTI, MOD -UV consistently outperforms baseline on all AP metrics except being comparable for APwith prior arts with ground-truth filtering. Specially, MOD -UV improves upon HASSOD∗(with ground truth filtering) on Box AP by 1.2 on nuScenes and 2.6 on KITTI, with notable improvements on AP by over 1.8 on nuScenes and 2.6 on KITTI. Even when compared on COCO which is the training set for HASSOD, MOD -UV outperforms prior arts on all metrics without ground-truth filtering except AP.

### v3

# MOD-UV: Learning Mobile Object Detectors from Unlabeled Videos

@@ Abstract
Embodied agents must detect and localize objects of interest, e.g. traffic participants for self-driving cars. Supervision in the form of bounding boxes for this task is extremely expensive. As such, prior work has looked at unsupervised instance detection and segmentation, but in the absence of annotated boxes, it is unclear how pixels must be grouped into objects and which objects are of interest. This results in over-/under-segmentation and irrelevant objects. Inspired by human visual system and practical applications, we posit that the key missing cue for unsupervised detection is motion: objects of interest are typically mobile objects that frequently move and their motions can specify separate instances. In this paper, we propose MOD-UV, a Mobile Object Detector learned from Unlabeled Videos only. We begin with instance pseudo-labels derived from motion segmentation, but introduce a novel training paradigm to progressively discover small objects and static-but-mobile objects that are missed by motion segmentation. As a result, though only learned from unlabeled videos, MOD-UV can detect and segment mobile objects from a single static image. Empirically, we achieve state-of-the-art performance in unsupervised mobile object detection on Waymo Open, nuScenes, and KITTI Datasets without using any external data or supervised models. Code is available at github.com/YihongSun/MOD-UV.
@@ 1 Introduction
Embodied agents such as self-driving cars must detect and localize objects of interest such as traffic participants to operate safely and effectively. Today, building such a detector requires the expensive and laborious annotation of millions of boxes over thousands of images. This process is so expensive that the largest detection dataset is orders of magnitude smaller than classification datasets and has much fewer classes. The limited set of classes further runs the risk of missing important object categories (e.g. snowplows for self-driving applications).
@@
These concerns have motivated research into unsupervised object detection techniques that automatically discover objects from unlabeled data [54, 55, 6]. Under the hood, these techniques use self-supervised features to segment unlabeled images and produce candidate object annotations which are then used to train a detector. However, while promising, these approaches often produce many uninteresting and irrelevant “objects” in cluttered scenes (e.g. buildings and roads) and over- or under-segment objects of interest (e.g. multiple detections partitioning a large bus or a single detection grouping a row of parked cars). These failures shouldn’t be surprising: after all, how can a completely unsupervised feature representation encode which assortment of windows, doors and wheels belongs together as an object, and which objects are of interest?
In this paper, we argue that a key missing cue for addressing the aforementioned issues in unsupervised instance detection is motion. In a practical sense, objects of interest are commonly mobile objects that frequently move. For example, robots performing navigation tasks must plan their trajectories carefully around objects that can move of their own volition. Thus, we argue that if we see similar groups of pixels frequently move of their own volition in unlabeled videos (e.g. vehicles and pedestrians in driving videos), this is sufficient information for building a mobile object detector that can detect such instances in static frames.
The importance of motion as a perceptual cue (the Gestalt principle of common fate) is well known [35]. Indeed, motion-based grouping is one of the first forms of grouping to appear developmentally in human infants [45] and can bootstrap other grouping cues [34]. There is also some prior work on using motion-based grouping in computer vision to produce (pseudo) ground-truth for feature learning [36] and discovering isolated, salient objects [11, 10]. However, there is a big-gap between motion segmentation and the kind of ground-truth we need for training a full-fledged instance-level object detector. First, motion segmentation produces a binary segmentation; this must be resolved into individual instances. Second, it only identifies moving objects and does not include objects (e.g. parked cars) that are static but mobile. Finally, motion segmentation only identifies nearby objects, since the pixel motion of faraway objects is too subtle to discern. Thus motion segmentation alone will still under-segment and miss many mobile objects: a problem for building object detectors.
@@
Here we propose a new training scheme to address these challenges. Our approach (MOD-UV, a Mobile Object Detector learned from Unlabeled Videos only; Figure 1) trains on unlabeled videos alone and produces a mobile object detector that can run on static frames. We first generate pseudo training labels from motion segmentation estimated by our prior unsupervised framework Dynamo-Depth [47]. We then propose a new training scheme to address the challenges above, resulting in a final mobile object detector that detects more mobile objects than the initial motion segmentation.
@@
We argue that motion as a cue is sufficient for unsupervised training of instance-level object detectors.
@@
We propose a new training scheme that trains on unlabeled videos to produce a mobile object detector that can run on static images.
@@ 2 Related Work
Learning to identify and localize objects from unlabeled images is a challenging task, since object information must be obtained without any explicit human annotations. A long line of work seeks to discover prominent objects in large image collections [9, 51, 52, 53]. However, these approaches are fundamentally limited by the quality of the object proposals. More recently, Locatello et al. [30] and DINOSAUR [38] consider object discovery as object-centric learning [18] and decompose a complex scene into independent objects. Nevertheless, the reconstruction objective is difficult to scale and likely to discover irrelevant patches as well.
More recent work has relied on the fact that bottom-up segmentation algorithms when applied to self-supervised pretrained representations yield good object proposals. Specifically, pseudo mask labels can be generated from DINO [7] features to train downstream object detectors [41, 42, 54]. MaskDistill [50] extends upon this by distilling from affinity graph produced by DINO [7] features, while TokenCut [56] and CutLER [55] use Normalized Cuts [39]. HASSOD [6] leverages hierarchical adaptive clustering, which improves the detection of small objects and object parts. This line of work now produces detectors that can run on static images, similar to our work. However, the detected objects can often be irrelevant (e.g. buildings and road) or over-/under-segment objects of interest. In contrast, our proposed MOD-UV discovers and detects a more meaningful and practical set of mobile objects instead, and can learn from their apparent motion in unlabeled videos only without relying on any additional datasets.
@@
In addition to unlabeled images, 3D information is also useful for discovering objects. Herbst et al. [21] and MODEST [62] discover non-persistent objects via multiple traversals with 3D sensors. Garcia et al. [16] discovers salient objects by late-fusing color and depth segmentation from RGB-D inputs, while Tian et al. [49] generates candidate segments from LiDAR 3D point clouds. In comparison, MOD-UV does not require any additional sensors or modalities beyond unlabeled videos, which allows our method to work in more general settings.
2.0.3 Unsupervised Object Detection/Discovery from Videos.
Inspired by Gestalt principle of common fate [35], another class of related work discovers objects via their apparent motions observed in videos [59, 61, 31]. By leveraging optical flow information from an input video, a binary segmentation of the moving objects can be extracted [36, 25, 57, 26, 60, 63, 64, 43]. Lian et al. [28] proposes further improvements for cases of articulated/deformable objects and shadow/reflections by relaxing the common fate assumption. Another line of work uses a reconstruction objective to identify the moving object [48, 1]. Du et al. [13] models explicit object geometry and physical dynamics by exploiting motion cues. Bao et al. [2] improves training for object-centric representation via an additional motion segmentation regularization, while SAVi++ [14] incorporates LiDAR data when training an object-centric video model. Unlike MOD-UV, these approaches do not build a static image detector. However, the output segmentation can be used as an initialization for our approach.
Closer to our work, Pathak et al. [36], Croitoru et al. [11] and Choudhury et al. [10] train a single-frame binary segmentation network on video frames as input and leverage object motion as supervision. Furthermore, LOCATE [44] applies graph-cut to obtain binary motion mask from DINO [7] and optical flow feature similarities, which in turn is treated as pseudo-labels for bootstrapped self-training of a downstream segmentation network. However, these techniques can only detect a single salient object per frame. In contrast, MOD-UV generalizes to multi-object detection beyond single-object saliency detection.
@@ 3 Method
3.0.1 Problem setup:
We assume an uncurated collection of unlabeled videos as input. In particular, we assume that these videos are obtained by an embodied agent observing, and optionally acting in the world. Solely from the unlabeled videos, the goal is to learn a detector that operates from a single frame and can detect and segment all mobile objects that can move of their own volition.
@@ 3.1 Initialization with Unsupervised Motion Segmentation
A key insight in MOD-UV is that if an object can move, it is likely that it does move many times in the collected data. Thus, we start by identifying moving objects in the videos; they can be initial seeds for learning about mobile objects.
@@
On the left of Figure 2, the initial pseudo-labels , while having high precision, fail to capture the large static objects, e.g. the black sedan in the bottom. However, a parked black sedan looks the same as a moving black sedan if all one has is a single frame. In other words, moving and static objects are indistinguishable when observed from a single frame.
@@
Also, since came from a heavily-regularized detector, we maintain and finetune the pseudo-labels for larger objects by training a second detector from scratch in parallel, on the training pair without down-scaling or padding. Notably, this is different from traditional scale jittering, since the singular detector would be discouraged from detecting small objects at larger scales due to the limitations in . Upon convergence, we have a Large-object detector trained at original scale and another Small-object detector trained at a reduced scale. After aggregating their predictions and resolving conflicting proposals, we obtain the final pseudo-labels, . We note that separating out large and small object detectors in this way has been explored in supervised face detection [22].
@@ 3.3 Implementation details
We follow the official code release by Dynamo-Depth[47] and train the system on Waymo Open [46]. During initial pseudo-label generation, we binarize the estimated motion mask via a threshold of and cluster the pseudo 3D points via DBSCAN [15] using a 10-by-10 local pixel neighborhood connectivity.
@@
During Moving2Mobile, we train a detector for 3 epochs, with scale jittering from to . Since early-stopping is applied, we adopt a lower confidence threshold of to compute the next round pseudo-labels . During Large2Small, we train both the large and small detectors for 20 epochs, with fixed scaling at and , respectively. As both are trained to convergence, we adopt a higher confidence threshold of and , respectively, to compute the next round pseudo-labels with aggregation. For Final round, we train the detector from scratch for 20 epochs, with scale jittering from to . The self-training in MOD-UV takes 27 hours on 1 NVIDIA A6000 GPU.
@@ 4.1 Experimental Setup
Datasets. For evaluation, we focus our attention to self-driving datasets since uncurated, unlabeled video data is available and detecting mobile objects is of interest for autonomous vehicles. We train both MOD-UV‡ and MOD-UV on Waymo [46] and compare our performance with baselines on Waymo. We then also evaluate generalization to nuScenes [4], KITTI [17], and COCO [29].
We split the 798 sequences from Waymo train set into 762 for training and 36 for validation. After method development concludes, we evaluate on the held-out 1,881 test images (averaging 28.4 mobile instances per image) from Waymo val set [33]. Additionally, nuScenes, KITTI, and COCO are only used for evaluating generalization. We test on 3,249 front-camera images (average of 8.2 mobile instances per image) from the nuImage validation set for nuScene and 7,481 images (average of 6.9 mobile instances per image) in the 2D Detection training set for KITTI. For COCO, we evaluate on 870 images (average of 3.8 mobile instances per image) in COCO val 2017 that contain ground vehicles.
Baselines. For the task of unsupervised mobile object detection, there is no directly comparable baselines to the best of our knowledge. Therefore, we consider methods for unsupervised object detection, namely CutLER [55] and HASSOD [6], as the closest points of comparison.
CutLER [55] uses normalized cuts on DINO features (trained on ImageNet) to generate pseudo labels that are used to train a Cascade Mask R-CNN [5] with ResNet-50 backbone. We evaluate the official checkpoint. Furthermore, we consider an additional baseline where is directly predicted via CutLER, which effectively ablates the use of motion cues, as denoted by CutLERL2S.
HASSOD [6] is a follow-up to CutLER. It discovers objects on COCO [29] via a hierarchical adaptive clustering of DINO features (trained from ImageNet). The hierarchical clustering yields three “levels”: objects, object parts, and object sub-parts. For consistency, we consider all three hierarchical levels for evaluation. As before, we use its official released checkpoint. Since MOD-UV‡ is trained on Waymo Open [46], we also consider a version of HASSOD solely trained on Waymo, which we denote as HASSOD†.
CutLER[55] and HASSOD[6] are trained to detect all objects in an image regardless of their ability to move. However, our evaluation and approach is focused on mobile objects. We therefore also consider an “oracle” version of these baselines where we additionally remove any CutLER and HASSOD predictions that overlap by less than 0.1 in IoU with the ground truth instances. These oracles, namely CutLER∗, CutLERL2S∗, HASSOD∗, and HASSOD†∗, are grayed out in the tables to indicate additional ground-truth-based filtering.
@@
Metrics. We evaluate both Average Recall (AR and AR) and Average Precision (AP and AP). Since the task is unsupervised and no semantic information is given during training, we follow prior arts [55, 6] and evaluate class-agnostic AR and AP by treating all predicted and ground truth instances as a single class of “foreground” or “mobile” objects.
@@ 4.2 Unsupervised Mobile Object Detection and Segmentation
Recall. We report AR0.5 (average recall with IoU), AR, AR, AR, and AR for both box and mask predictions. MOD-UV significantly outperforms prior arts across all recall metrics except the recall for large objects where it is comparable. The improvement is especially large for small objects (4.7 higher AR than the nearest competitor). Compared to a supervised Mask R-CNN trained on COCO [29], MOD-UV closes the gap in Box AR from 11.6 to 4.8. Our gains are also much larger on the AR0.5 metric (nearly 2 prior state-of-the-art). This suggests that we detect significantly more objects than prior work, but their localization can be improved. Even so, we still show a -point improvement on overall AR. We also found HASSOD to underperform when trained solely on Waymo, which we suspect is due to the uncurated nature of self-driving scenes.
Precision. We report AP at an overlap threshold of 0.5, as well as AP, AP, AP and AP. Since CutLER and HASSOD are trained to detect all objects in an image regardless of their ability to move, we also compare to oracle versions of these techniques with ground-truth-based filtering. On Waymo, MOD-UV significantly outperforms prior arts across all precision metrics. Even with ground-truth filtering (in gray), MOD-UV still consistently outperforms baselines on all precision metrics (except for larger objects where it is comparable). Specifically, MOD-UV outperforms the nearest competitor (with ground-truth filtering) by 4.1 on Box AP and 3.4 on Mask AP, and is 4.5 higher on AP. Intriguingly, compared to a supervised Mask R-CNN trained on COCO [29], MOD-UV closes the gap in Mask AP from 4.3 to 0.8 points.
@@
Recall. As shown in Table 2 and Table 3, on nuScenes and KITTI, MOD-UV consistently outperforms prior arts across all AR metrics except for large objects, achieving a more than 1.5 improvement on AR0.5 over the nearest competitor on nuScenes. MOD-UV also shows large gains on small objects, improving AR by 2.4 on nuScenes and over 1.7 on KITTI.
Finally, we evaluate on COCO, which is in-domain for HASSOD and a big domain shift for MOD-UV. Notably, MOD-UV maintains superiority on AR and AR at IoU, while being comparable to HASSOD on AR and AR.
Precision. This improvement is also seen in AP. On both nuScenes and KITTI, MOD-UV consistently outperforms baseline on all AP metrics except being comparable for APwith prior arts with ground-truth filtering. Specially, MOD-UV improves upon HASSOD∗(with ground truth filtering) on Box AP by 1.2 on nuScenes and 2.6 on KITTI, with notable improvements on AP by over 1.8 on nuScenes and 2.6 on KITTI. Even on COCO, MOD-UV outperforms prior arts on all metrics without ground-truth filtering except AP.